OpenStudy (anonymous):

-3(4t)+5t-6 show work

OpenStudy (saifoo.khan):

-12t + 5t - 6

OpenStudy (saifoo.khan):

-7t - 6

OpenStudy (anonymous):

khan what the heck

OpenStudy (saifoo.khan):

LoL.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

-3(4t)+5t-6 -12t+5t-6 -7t-6 and if you meant -3(4t)+5t-6 = 0 -12t+5t-6 = 0 -7t-6 = 0 7t = -6 t = -6/7

OpenStudy (saifoo.khan):

ant, u r wrong.

OpenStudy (saifoo.khan):

sant*

OpenStudy (saifoo.khan):

this isnt an equation. this is simple expression

OpenStudy (anonymous):

I showed the same result as you and even the root of it

OpenStudy (anonymous):

well it is the root then

OpenStudy (saifoo.khan):

u cannot solve it further.

OpenStudy (anonymous):

yes, what I mean is Result: -7t - 6 Root: t = -6/7

OpenStudy (saifoo.khan):

still innc.

OpenStudy (anonymous):

then you must be right

OpenStudy (anonymous):

so what is the answer?

OpenStudy (saifoo.khan):

Mine.

OpenStudy (anonymous):

lol k

OpenStudy (anonymous):

-7 - 6?

OpenStudy (saifoo.khan):

-7t - 6

OpenStudy (anonymous):

8y- (2y+3)-19

OpenStudy (anonymous):

Will you help with a couple more?

OpenStudy (anonymous):

I can help

OpenStudy (anonymous):

K

OpenStudy (anonymous):

6 y-22

OpenStudy (saifoo.khan):

sorry, rose i really gotta go now!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

-7y=21

OpenStudy (anonymous):

y = -3

OpenStudy (anonymous):

z-23=-28

OpenStudy (anonymous):

z=-5